# The Rhythm of Aging Stability and Drift in Human Senescence

### v1

@@ 1 Introduction
Aging is the gradual decline in physiological function — what we see as graying hair, slower steps, and growing vulnerability to illness and injury. Beneath these visible signs lies senescence, the biological process that drives aging (Comfort, , 1964). Demographers focus on actuarial senescence — the age-related rise in mortality risk — which, in most adult populations, follows an exponential curve.
@@
If the pace of aging is fixed, then aging is not speeding up or slowing down — it’s just being postponed (Vaupel, , 2010). But if the rate of aging is truly changing, then something more fundamental is happening (Kirkwood and Austad, , 2000). It would mean that the biology of aging itself is evolving — or being altered by the environment, behavior, or historical events (Finch and Crimmins, , 2004; Crimmins and Beltrán-Sánchez, , 2011).
@@
This matters because when we estimate $b$ cohort by cohort, we assume we are tracing a biological process. But we may be picking up the long echo of shared historical event — nudging the estimated rate of aging up or down, year by year, in ways that mimic biological drift. Over time, these nudges can pile up, creating what looks like a change in the slope of mortality — even if aging itself hasn’t changed at all (Zarulli et al., , 2012; Salinari and De Santis, , 2014; Horiuchi and Wilmoth, , 1998).
These kinds of latent effects are hard to see directly. But their signature is familiar: they accumulate gradually, move in one direction for a while, then turn. We can think of this as a stochastic process—more specifically, a random walk (e.g., Grimmett and Stirzaker, , 2020; Hamilton, , 2020). If period-driven shocks behave this way, they could generate the illusion of a shifting $b$ — even when the underlying biology stays constant (Yashin et al., , 2000).
@@ 2 Background and Motivation
Estimates of the rate of aging, $b$, often vary across birth cohorts. Sometimes these values drift gently up or down; other times they fluctuate without clear direction. Many of these differences are statistically significant, but their biological interpretation remains uncertain. Do these fluctuations reflect real variation in how humans age? Or are they shaped by model assumptions, measurement noise, or shared historical exposures?
@@ 3.3 Decomposing Variation in $b$
When we estimate $b$ cohort by cohort, we observe subtle but persistent fluctuations. These movements are too structured to be dismissed as noise — yet they don’t follow a clear trend. Figure 1 illustrates this drift, alongside its first differences and a possible cumulative component.
@@
This pattern raises a central question: what if these changes aren’t biological at all? Period events — such as wars, pandemics, or economic disruptions — affect many cohorts at once, just at different ages. If their effects accumulate over time, they may leave a cohort-shaped fingerprint in the estimates. What appears to be a change in $b$ might simply be the echo of shared historical experience — not a shift in the biology of aging.
@@ 3.4 Estimation Process
We implemented the model in a Bayesian framework using Stan (Stan Development Team, 2025b, ), via its R interface RStan (Stan Development Team, 2025a, ). Four independent Markov chains were run for 6,000 iterations each, with the first 4,000 iterations discarded as warm-up and the remaining 8,000 retained as posterior samples. Convergence was assessed using the $\hat{R}$ diagnostic, which remained below 1.05 for all parameters (Vehtari et al., , 2021), indicating reliable mixing and convergence.
Point estimates reported are the maximum a posteriori (MAP) values, with uncertainty summarized using 95% Highest Posterior Density (HPD) intervals (Patricio and Missov, , 2023).
@@ 4.1 Observed Variation in the Rate of Aging
To further explore the structure of variation in $\log b_{t}$, we plot the first differences across cohorts (Figure 3). These differences cluster closely around zero, with no indication of sustained directional change. The smooth, mean-reverting nature of these fluctuations supports the idea that what we observe is not biological shift, but a slow drift consistent with a latent cumulative process — such as a random walk. This interpretation is further strengthened by formal stationarity tests (ADF, KPSS, Phillips-Perron for $\alpha=0.05$), which confirm that the log-differences are stationary, suggesting $\log b_{t}$ behaves like an integrated process (e.g., Hamilton, , 2020).
@@ 4.3 What Remains After the Drift?
What’s left are the residuals: small, irregular fluctuations shown in Figure 6. They pass standard stationarity tests (ADF, KPSS, Phillips-Perron for $\alpha=0.05$), and they carry no visible signal. In other words, once we account for shared shocks, what remains is consistent with random noise.
@@ 5 Discussion
This study set out to revisit Vaupel’s hypothesis: that the rate of aging, $b$, is a biological constant. We asked whether the fluctuations in $b$ observed across cohorts reflect real shifts in the pace of senescence — or whether they arise from something else: the cumulative effect of shared historical shocks.
@@
The results strongly support the latter. Once we isolate senescent mortality and decompose the variation in $b_{t}$, we find no statistical trend. The apparent shifts in the rate of aging are not consistent with a changing biology. Instead, they trace the signature of a latent process that builds slowly over time — a cumulative period effect. In this view, aging itself remains stable. What moves is the historical context.
This reinforces what Vaupel originally proposed: that aging is not slowing down, only starting later. The slope stays the same — it’s just been pushed forward. Our findings align with earlier studies that found limited variation in $b$ (e.g., Barbi et al., , 2003), but offer a new explanation for why those variations appear in the first place. Where previous research saw fluctuation, we show structure.
@@
At the same time, the evidence is not definitive. The confidence intervals around $b_{t}$ — particularly for small population countries — remain wide. The results are statistically consistent with stability, but more precise estimates would require additional data: broader coverage, more cohorts, or longer time series. These would allow for tighter inference and a clearer distinction between signal and noise.
There are also modeling assumptions that should be noted. The latent period effect is structured as a random walk, which captures gradual accumulation but may overlook abrupt shocks or nonlinear interactions. The decomposition is linear, and we treat frailty as stable across cohorts — assumptions that may not hold in all settings. Future work could test alternative specifications, introduce flexible latent processes, or incorporate Bayesian structural time series approaches.
Finally, while this study focuses on national-level mortality, the approach could be extended to explore heterogeneity across subgroups. Do historical shocks leave different fingerprints by sex, education, or socioeconomic status? Does the filtering of senescence vary by social context? These are open questions that matter — not just for understanding aging, but for anticipating how populations will age in the future.
@@
This is, to our knowledge, the first study to explicitly decompose the cohort-level variation in $b$ into a biological constant, a trend, and a latent period effect. In doing so, we offer a framework for biodemographic filtering — a way to separate the signal of aging from the noise of period events. The model is simple, extensible, and grounded in demographic logic. As interest in senescence deepens, we hope this structure helps clarify what, exactly, is changing — and what is not.
@@ 6 Conclusion
The Gompertz slope, $b$, is one of the most widely used measures of the pace of aging. Yet its apparent variation across cohorts has raised persistent doubts about whether it is truly stable — or quietly changing.
This study shows that most of the fluctuation in $b$ is not biological. Once we filter out non-senescent mortality and account for the accumulation of shared period effects, $b$ becomes strikingly consistent. There is no meaningful trend, and no evidence that the biology of aging is shifting across cohorts.
These findings provide new support for Vaupel’s hypothesis: that aging proceeds at a fixed biological pace. What has changed over time is not how fast people age, but when aging begins. The rhythm stays the same — it simply starts later.

### v2

@@ 1 Introduction
Aging is the gradual decline in physiological function — what we see as graying hair, slower steps, and growing vulnerability to illness and injury. Beneath these visible signs lies senescence, the biological process that drives aging (Comfort, , 1964). Demographers focus on actuarial senescence — the age-related rise in mortality risk — which, in most adult populations, follows an exponential curve (Strehler and Mildvan, , 1960; Olshansky et al., , 1990).
@@
If the pace of aging is fixed, then aging is not speeding up or slowing down — it’s just being postponed (Vaupel, , 2010). But if the rate of aging is truly changing, then something more fundamental is happening (Kirkwood and Austad, , 2000). It would mean that the biology of aging itself is evolving — or being altered by the environment, behavior, or historical events (Finch and Crimmins, , 2004; Crimmins and Beltrán-Sánchez, , 2011; Olshansky et al., , 1990).
@@
This matters because when we estimate $b$ cohort by cohort, we assume we are tracing a biological process. But we may be picking up the long echo of a shared historical event — nudging the estimated rate of aging up or down, year by year, in ways that mimic biological drift. Over time, these nudges can pile up, creating what looks like a change in the slope of mortality — even if aging itself has not changed at all (Zarulli et al., , 2012; Salinari and De Santis, , 2014; Horiuchi and Wilmoth, , 1998).
These kinds of latent effects are hard to see directly. But their signature is familiar: they accumulate gradually, move in one direction for a while, then turn. Statistically, they resemble a stochastic process — more precisely, a random walk (e.g., Grimmett and Stirzaker, , 2020; Hamilton, , 2020). If period-driven shocks follow this pattern, they could mimic a changing $b$, even when the biology of aging holds steady (Yashin et al., , 2000). As Alter and Riley, (1989) noted, trajectories of frailty and mortality are often shaped not just by individual biology, but by shared historical conditions.
@@ 2 Background and Motivation
Estimates of the rate of aging, $b$, often vary across birth cohorts. Sometimes these values drift gently up or down; other times they fluctuate without clear direction. Many of these differences are statistically significant, but their biological interpretation remains uncertain. Earlier studies raised a similar concern: that what looks like change in aging may reflect measurement artifacts, not biological processes (Strehler and Mildvan, , 1960; Olshansky et al., , 1990). Do these fluctuations reflect real variation in how humans age? Or are they shaped by model assumptions, statistical noise, or shared historical exposures?
@@ 3.3 Decomposing Variation in $b$
When we estimate $b$ cohort by cohort, we observe subtle but persistent fluctuations. These movements are too structured to be dismissed as noise — yet they do not follow a clear trend. Figure 1 illustrates this drift, alongside its first differences and a possible cumulative component.
@@
This pattern raises a central question: what if these changes are not biological at all? Period events — such as wars, pandemics, or economic disruptions — affect many cohorts at once, just at different ages. If their effects accumulate over time, they may leave a cohort-shaped fingerprint in the estimates. What appears to be a change in $b$ might simply be the echo of shared historical experience — not a shift in the biology of aging.
@@ 3.4 Estimation Process
Before estimating the model, we center the time variable $t$ by subtracting the mean cohort year. This ensures that the intercept in Equation 3 corresponds to the average cohort, rather than an arbitrary reference year. Centering reduces correlation between the intercept and slope, leading to more stable estimates and improved mixing in the Bayesian sampler. [^1] Centering time is a standard practice in time-series and regression models to reduce correlation between parameters (Hamilton, , 2020; Gelman and Hill, , 2007). In our setup, it also helps ensure that small cohort-to-cohort shifts are captured by the trend term ($\beta t$), rather than being absorbed into the latent drift process.
We implemented the model in a Bayesian framework using Stan (Stan Development Team, 2025b, ), accessed via its R interface RStan (Stan Development Team, 2025a, ). We ran four independent Markov chains with 6,000 iterations each, discarding the first 4,000 as warm-up and retaining the remaining 2,000 as posterior samples from each chain — yielding a total of 8,000 posterior samples. Convergence was assessed using the $\hat{R}$ diagnostic, which remained below 1.05 for all parameters (Vehtari et al., , 2021), indicating reliable mixing and convergence.
Point estimates reported are maximum a posteriori (MAP) values, and uncertainty is summarized using 95% Highest Posterior Density (HPD) intervals (Patricio and Missov, , 2023).
@@ 4.1 Observed Variation in the Rate of Aging
To further explore the structure of variation in $\log b_{t}$, we plot the first differences across cohorts (Figure 3). These differences cluster closely around zero, with no indication of sustained directional change. The smooth, mean-reverting nature of these fluctuations supports the idea that what we observe is not biological shift, but a slow drift consistent with a latent cumulative process — such as a random walk. This interpretation is further supported by formal stationarity tests, [^2] We apply Augmented Dickey-Fuller (ADF), Kwiatkowski-Phillips-Schmidt-Shin (KPSS), and Phillips-Perron (PP) tests at the 5% significance level. These tests differ in their null hypotheses and sensitivity to autocorrelation and heteroskedasticity — ADF and PP test for a unit root, while KPSS tests for stationarity — which provides a more robust check for stationarity. which confirm that the log-differences of $\log b_{t}$ are stationary. This supports our treatment of $b_{t}$ as a non-stationary but mean-reverting process — consistent with a latent random walk (e.g., Hamilton, , 2020).
@@ 4.3 What Remains After the Drift?
What’s left are the residuals: small, irregular fluctuations shown in Figure 6. They pass standard stationarity tests, [^3] Stationarity was confirmed using the ADF, KPSS, and PP tests, all at the 5% significance level. As before, these tests complement one another by probing different null hypotheses and responding differently to autocorrelation and heteroskedasticity. and they carry no visible signal. In other words, once we account for shared shocks, what remains is consistent with random noise.
@@ 5 Discussion
This study set out to revisit Vaupel’s hypothesis: that the rate of aging, $b$, is biologically constant. We asked whether fluctuations in $b$ reflect real shifts in the biology of senescence — or the echo of shared historical shocks.
@@
The results strongly support the latter. Once we isolate senescent mortality and decompose the variation in $b_{t}$, we find no meaningful trend. What looks like a change in the pace of aging turns out to be drift: a latent process that accumulates period effects over time. In this view, senescence remains stable. What varies is the environment around it.
Vaupel’s hypothesis gains new support from this decomposition. The underlying biological signal of $b$ aligns with his original claim (Vaupel, , 2010), while the “noise” others observed is largely explained by historical shocks. Where Barbi et al., (2003) saw model sensitivity, we find structured drift. And while Zarulli et al., (2012) and Salinari and De Santis, (2014) interpreted cohort fluctuations as potential biological change, our results suggest those shifts reflect shared exposures in calendar time. As Alter and Riley, (1989) noted, mortality is deeply embedded in historical context — what appears biological may in fact be social.
The stability of the Gompertz slope also echoes biological theories that emphasize conserved mechanisms of aging. The “hallmarks of aging” framework (López-Otín et al., , 2013, 2023) highlights core processes — like telomere attrition and genomic instability — that operate consistently across species. Our findings suggest these same mechanisms may remain stable across time, even as mortality patterns shift.
If the pace of aging does not change, but its onset does, then the question becomes: what governs the timing of senescence? Is it early-life conditions, cumulative exposures, or broader social trajectories? Like Finch and Crimmins, (2004) and Crimmins and Beltrán-Sánchez, (2011), we find that aging’s timing may respond to environmental forces, even if its tempo does not. And the widening gaps in healthy life expectancy (Crimmins and Saito, , 2001) underscore the unequal ways in which this timing plays out.
This distinction between when aging begins and how fast it proceeds touches core debates in biodemography and gerontology. It bears on whether healthspan can be extended or merely redistributed, whether longevity gains compress morbidity or stretch mortality. It suggests the future of aging lies not in slowing the clock, but in delaying when it starts to tick.
@@
Still, these findings are not definitive. The confidence intervals around $b$ — especially in smaller populations — remain wide. The results are statistically consistent with stability, but more precise estimates would require additional data: broader coverage, more cohorts, longer time series, and richer subgroup analysis.
Methodologically, the model could also be extended — not to undermine the hypothesis of stability, but to test it more rigorously. The latent period effect here is modeled as a random walk without drift. But one might relax that assumption. A drift term, or an autoregressive structure, could test whether mortality patterns return to baseline or nudge persistently in one direction. If the drift is negligible, it would further support the idea that aging’s pace is stable — buffeted by shocks, but not reoriented by them.
This framework also opens new directions. While we focus here on national populations, the same approach could be applied to subgroups — by sex, education, income, or region. Do all groups experience aging as a stable rhythm? Or do life course inequalities leave lasting marks on senescence? These are not just statistical questions, but fundamental demographic ones.
@@
To our knowledge, this is the first study to decompose the variation in $b$ into a biological constant, a trend, and a latent period effect. In doing so, we offer a way to filter the signal of aging from the noise of period events. The model is simple, extensible, and grounded in demographic logic. As interest in senescence deepens, we hope it helps clarify what, exactly, is changing — and what is not.
@@ 6 Conclusion
The Gompertz slope, $b$, remains one of the most widely used measures of the pace of aging. Yet its modest variation across cohorts has long raised a deeper question: is the rate of aging truly stable — or quietly shifting?
This study suggests it is stable. Once we filter out non-senescent mortality and account for the accumulation of shared period effects, $b$ becomes strikingly consistent. There is no meaningful trend, and no evidence that the biology of aging is shifting across cohorts. The variation that remains is not biological drift, but the footprint of history — the cumulative result of shocks experienced in calendar time.
These findings provide new support for Vaupel’s hypothesis: that aging proceeds at a fixed biological pace. What has changed is not how fast people age, but when aging begins. The rhythm stays the same — it simply starts later.